I have this math problem I've​ been struggling with, it has multiple parts but it's basically the same idea. the question is solve the problems. With each step, please.Find the real solutions of each equation. (x + 2)4/3-5(x + 2)2/3 + 6-0 3x3 + 4x2-27x + 36
Answer #1

Solution: 1)

  \left ( x+2 \right )^{}\frac{4}{3}-5\left ( x+2 \right )^{\frac{2}{3}}+6=0

  \therefore[ \left ( x+2 \right )^{2}]^{}\frac{2}{3}-5\left ( x+2 \right )^{\frac{2}{3}}+6=0

put \left ( x+2 \right )^{\frac{2}{3}}=u

\therefore u^{^{2}}-5u+6=0

\therefore \left ( u-3 \right )\left ( u-2 \right )=0

\therefore u=3\: \: or\: \: \: u=2

Resubstitute the value of \left ( x+2 \right )^{\frac{2}{3}}

\therefore \left ( x+2 \right )^{\frac{2}{3}}=3\: \: \: or\: \: \left ( x+2 \right )^{\frac{2}{3}}=2

\therefore x+2=3^{\frac{3}{2}}\: \: \: or\: \: \: \: x+2=2^{\frac{3}{2}}

\therefore x=3^{\frac{3}{2}}-2\: \: \: or\: \: \: \: x=2^{\frac{3}{2}}-2

2)   3x^{3}+4x^{2}-27x-36=0

By synthetic division

3 3 4 -27 -36
9 39 36
3 13 12 0

\therefore \left ( x-3 \right )\left ( 3x^{2} \right +13x+12)=0

\therefore \left ( x-3 \right )\left ( x+3 \right )\left ( 3x+4 \right )=0

\therefore \left ( x^{2}-9 \right )\left ( 3x+4 \right )=0

\therefore x=\pm 3\: \: \: or\: \: \: x=-\frac{4}{3}

3) 2x-1=-\sqrt{2-x}

squaring on both side we get

\left ( 2x-1 \right )^{2}=2-x

\therefore 4x^{2}-4x+1=2-x

\therefore 4x^{2}-3x-1=0

\therefore \left ( x-1)\left ( 4x+1 \right ) \right =0

\therefore x=1\: \: \: or\: \: \: x=-\frac{1}{4}

4)\sqrt{2x-5}-\sqrt{x-3}=1

Sqquaring on both side

\therefore 2x-5+x-3-2\sqrt{2x-5}\sqrt{x-3}=1

\therefore 3x-9=2\sqrt{2x-5}\sqrt{x-3}

Squaring on both side

\therefore 9x^{2}-54x+81=8x^{2}-44x+60

\therefore x^{2}-10x+21=0

\therefore \left ( x-7 \right )\left ( x-3 \right )=0

\therefore x=7\: \: \: or\: \: \: x=3

5)   \sqrt[4]{5x^{2}-6}=x

5x^{2}-6=x^{4}

\therefore x^{4}-5x^{2}+6=0

\therefore \left ( x^{2}-2 \right )\left ( x^{2} -3\right )=0

x=\pm \sqrt{2}\: \: \: or\: \: \: \: x=\pm \sqrt{3}